test bank

John Wall Inc. is launching a line of "2" branded items in a 2-year project that involves equipment that will be purchased today for $170000, relevant annual sales of $100000, relevant annual costs of $70000, and a tax rate of 20%. What is OCF expected to in 2nd year of the project if MACRS depreciation is used where the depreciation rates in years 1, 2, 3, and 4 are 50%, 30%, 20%, and 10%, respectively?

Selected Answer: 340000Correct Answer: 34,200
Answer #1

To calculate the Operating Cash Flow (OCF) in the 2nd year of the project, we need to determine the relevant cash flows for that year. Here's how we can do it step-by-step:

 

Calculate the depreciation expense for each year using the MACRS depreciation method.

 

Determine the relevant cash flows for each year, considering sales, costs, depreciation, and taxes.

 

Calculate the OCF for the 2nd year of the project.

 

Let's start with the calculations:

 

Step 1: Calculate the depreciation expense for each year using the MACRS depreciation method:

 

Year 1:

Depreciation expense = Equipment cost × MACRS depreciation rate for Year 1

Depreciation expense = $170,000 × 50% = $85,000

 

Year 2:

Depreciation expense = Equipment cost × MACRS depreciation rate for Year 2

Depreciation expense = $170,000 × 30% = $51,000

 

Step 2: Determine the relevant cash flows for each year:

 

Year 1:

Sales revenue = $100,000

Costs = $70,000

Depreciation = $85,000 (calculated above)

Taxable income = Sales - Costs - Depreciation

Taxable income = $100,000 - $70,000 - $85,000 = -$55,000 (negative because of the large depreciation expense)

 

Taxes = Tax rate × Taxable income

Taxes = 20% × -$55,000 = -$11,000 (negative because of the tax shield from the loss)

 

Net cash flow (NCF) = Sales - Costs - Taxes

NCF = $100,000 - $70,000 - (-$11,000) = $41,000

 

Year 2:

Sales revenue = $100,000

Costs = $70,000

Depreciation = $51,000 (calculated above)

Taxable income = Sales - Costs - Depreciation

Taxable income = $100,000 - $70,000 - $51,000 = -$21,000 (negative because of the depreciation expense)

 

Taxes = Tax rate × Taxable income

Taxes = 20% × -$21,000 = -$4,200 (negative because of the tax shield from the loss)

 

NCF = Sales - Costs - Taxes + Depreciation

NCF = $100,000 - $70,000 - (-$4,200) + $51,000 = $85,800

 

Step 3: Calculate the OCF in the 2nd year of the project:

 

OCF = Net cash flow (NCF) + Depreciation

OCF = $85,800 + $51,000 = $136,800

 

Therefore, the Operating Cash Flow (OCF) expected in the 2nd year of the project is $136,800.


answered by: anonymous